How much growth occurred in the state of North Dakota between 1998 and 2014 if real GDP (measured in

    millions of 2009 chained dollars) was $21,876 in 1998 and $48,233 in 2014?
Answer #1

The growth = % change in real GDP between 1998 and 2014 = (48,233 - 21,876)/21,876 * 100 = 26,357/21,876 * 100 = 1.2048 * 100 = 120.48%

So, the real GDP increased by 120.48%